A student dissolves 18. g of sucrose (C12H22On) in 150. mL of a solvent with a density of 1.09 gml. The student notices that the volume of the solvent does not change when the sucrose dissolves in it. Calculate the molarity and molality of the students solution. Be sure each of your answer entries has the correct number of significant digits. molarity = molality =
Answer #1

1)

Molar mass of C12H22O11,

MM = 12*MM(C) + 22*MM(H) + 11*MM(O)

= 12*12.01 + 22*1.008 + 11*16.0

= 342.296 g/mol

mass(C12H22O11)= 18 g

number of mol of C12H22O11,

n = mass of C12H22O11/molar mass of C12H22O11

=(18.0 g)/(342.296 g/mol)

= 5.259*10^-2 mol

volume , V = 150 mL

= 0.15 L

Molarity,

M = number of mol / volume in L

= 5.259*10^-2/0.15

= 0.35 M

Answer: 0.35 M

2)

m(solvent)= density * volume

= 1.09 g/mL * 150 mL

= 163.5 g

= 0.1635 Kg

Molality,

m = number of mol / mass of solvent in Kg

=(5.259*10^-2 mol)/(0.1635 Kg)

= 0.32 molal

Answer: 0.32 molal